OpenStudy (anonymous):

create a function h(x) that has a vertical asymptotes at x=3/5 and x=-5/2 and a horizontal asymptote of y=8

OpenStudy (anonymous):

One such function is \[ f(x) =\frac { a x^2}{ (c x +d)(mx + n)} \] Let us try to find a, b,c, m and n

OpenStudy (anonymous):

c x + d = 5 x - 3\\ mx + n= 2 x +5\\ a = 8 (m c) = 80\\ Why?